A bushed pin type flexible coupling is used to transmit 10 kW power at 720 rpm. The design torque is 150% of rated torque. The keys have square cross-section. The permissible stresses are: For shaft and key material, τ = 66.67 N/mm2 , σc = 200 N/mm2 ; For pin material, τ = 35 N/mm2 , σt = 133 N/mm2 ; For flange material τ = 16.67 N/mm2 . The permissible bearing pressure for rubber bushes is 1 N/mm2 . The number of bushes is 4. Design the bushed pin flexible coupling.                                                                                                                                         
Answer #1

Snfaitre N 720 6pm dugn benque1-ated Permissbu Atrssus tor shafts Vsy and 8S N/mm てこ 133-33 NImm 2 T B3NImm b h Diameto ofshanehub is teatre sudted to iongl JR(dd inder hellus a mamant -x(so 5t) =SIS242 -80mn 2 32 8-65N/mn as242 40) てく1663 Nim fangeAuma huahis Q198943 68) AN (lo)4) 3154 or 35 mm ndsing str u fhias actoung o ulatd by and Terques pew( P) fleing ro each On 94198993 681 (25) (3) (60) 3t8 nduced in h heAhiar and compruasive tansnan urmssibt kanmut ys ane witkig
Answer #2

A bushed pin type flexible coupling is used to transmit 10 kW power at 720 rpm. The design torque is 150% of rated torque. The keys have square cross-section. The permissible stresses are: For shaft and key material, τ = 66.67 N/mm2 , σc = 200 N/mm2 ; For pin material, τ = 35 N/mm2 , σt = 133 N/mm2 ; For flange material τ = 16.67 N/mm2 . The permissible bearing pressure for rubber bushes is 1 N/mm2 . The number of bushes is 4. Design the bushed pin flexible coupling. Explain Axle, Spindle, Counter shaft and line-shaft with their examples.